A spring-loaded toy gun is designed to fire a rubber ball by compressing it into a spring a total of 30 cm. When fired vertically upward, the ball leaves the gun with a speed of 2.6 m/s. With what speed (in m/s) will the ball leave the gun if it is fired horizontally instead?
Answer #1

using conservation of energy

0.5 k x^2 = mg x + 0.5 m v^2

0.5* k * 0.3^2 = m ( 9.8*0.3 + 0.5 * 2.6^2)

k/m = 140.44

now,

0.5 k x^2 = 0.5 m v^2

k/m x^2 = v^2

140.44 * 0.3^2 = v^2

v = 3.555 m/s

=====

Comment before rate in case any doubt, will reply for sure.. goodluck
Answer #2

using conservation of energy

0.5 k x^2 = mg x + 0.5 m v^2

0.5* k * 0.3^2 = m ( 9.8*0.3 + 0.5 * 2.6^2)

k/m = 140.44

now,

0.5 k x^2 = 0.5 m v^2

k/m x^2 = v^2

140.44 * 0.3^2 = v^2

v = 3.555 m/s

Comment in case any doubt please rate my answer....